find the values of sin cos tan 75

Grade:9

2 Answers

Ganga sagar Kothashiv
11 years ago

sin(75)=(31/2+1)/(2*21/2)

cos(75)=(31/2--1)/(2*21/2)

tan(75)=(31/2+1)/(31/2-1)

sheik arbaz
11 years ago

0.9

0.25

3.3